you have 13000 to invest in a stock portfolio. your choice are stock x with an expected return of 15 percent and stock y with an expected return of 9 percent. assume your goal is to create a portfolio with an expect return of 12.35
Answer #1
Let investment in X be "x"
Investment In Y = 13000 - x
RETURN FROM X + RETURN FROM Y = REQUIRED RETURN
x * 15% + (13000 - x) * 9% = 13000 * 12.35%
0.15x + 1170 - 0.09x = 1605.5
0.06x = 435.50
x = 7258.33
Investment in X = $ 7258.33
Investment In Y = 13000 - x = 13000 - 7258.33 = $ 5741.67